# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ 4 An Effective Algorithm
$\mathcal{L}_{fair}(\bm{\theta})=\big{|}\frac{1}{m}\sum_{(\mathbf{x}_{i},z_{i})\in\mathcal{B}}g(\hat{f}(\mathbf{x}_{i},\bm{\theta}),z_{i})\big{|}+\big{|}\frac{1}{m}\sum_{(\mathbf{x}_{j},z_{j})\in\mathcal{B}^{\prime}}g(\hat{f}(\mathbf{x}_{j},\bm{\theta}),z_{j})\big{|}$
10:         $\mathcal{L}(\bm{\theta})=\mathcal{L}_{cls}(\bm{\theta})+\lambda_{1}\cdot\mathcal{L}_{inv}(\bm{\theta})+\lambda_{2}\cdot\mathcal{L}_{fair}(\bm{\theta})$
11:         $\bm{\theta}\leftarrow\text{Adam}(\mathcal{L}(\bm{\theta}),\bm{\theta},\eta_{p})$
12:         $\lambda_{1}\leftarrow\text{max}\{[\lambda_{1}+\eta_{d}\cdot(\mathcal{L}_{inv}(\bm{\theta})-\gamma_{1})],0\}$, $\lambda_{2}\leftarrow\text{max}\{[\lambda_{2}+\eta_{d}\cdot(\mathcal{L}_{fair}(\bm{\theta})-\gamma_{2})],0\}$
@@ 5 Experiments
Settings. We evaluate the performance of our FDDG on four benchmarks. To highlight each domain $e$ and its fair dependence score $\rho^{e}$, we summarize the statistic in Tab. 1. Three image datasets ccMNIST, FairFace [21], YFCC100M-FDG [22], and one tabular dataset New York Stop-and-Frisk (NYSF) [23] are conducted on FDDG against 17 state-of-the-art baseline methods that fall into two categories, state-of-the-art DG methods (RandAug [^1] RandAug (or ColorJitter) is a naive built-in function in Torch used for image transformations. It randomly changes the brightness, contrast, saturation, and hue of given images., ERM [24], IRM [8], GDRO [25], Mixup [26], MLDG [27], CORAL [28], MMD [29], DANN [30], CDANN [31], DDG [13], and MBDG [12]) and fairness-aware methods in changing environments (DDG-FC, MBDG-FC  [^2] DDG-FC and MBDG-FC are two fairness-aware DG baselines that built upon DDG [13] and MBDG [12], respectively. These extensions involve the straightforward addition of fairness constraints defined in Defn. 1 to the loss functions of the original models., EIIL [4], FarconVAE [5], FATDM [6]). Three metrics are used for evaluation. Two of them are for fairness quantification, Demographic Parity (DP) [14] and the Area Under the ROC Curve (AUC) between predictions of sensitive subgroups [32]. Notice that the AUC metric is not the same as the one commonly used in classification based on TPR and FPR. The intuition behind this AUC is based on the nonparametric Mann-Whitney U test, in which a fair condition is defined as the probability of a randomly selected sample $\mathbf{x}_{-1}$ from one sensitive subgroup being greater than a randomly selected sample $\mathbf{x}_{1}$ from the other sensitive subgroup is equal to the probability of $\mathbf{x}_{1}$ being greater than $\mathbf{x}_{-1}$ [33, 34]. A value of DP closer to $1$ indicates fairness and $0.5$ of AUC represents zero bias effect on predictions. Due to space limits, we defer a detailed description of the experimental settings (including datasets, baselines, evaluation metrics, etc.) in Appendix C and complete results on all baselines and datasets in Appendix F.
@@
Dataset Domain, $e$ Sensitive label, $Z$ Class label, $Y$ ($e$, $\rho^{e}$), $\forall e\in\mathcal{E}$
@@
FairFace race gender age (B, 0.91), (E, 0.87), (I, 0.58), (L, 0.48), (M, 0.87), (S, 0.39), (W, 0.49)
YFCC100M-FDG year location indoor/outdoor ($d_{0}$, 0.73), ($d_{1}$, 0.84), ($d_{2}$, 0.72)
NYSF city race stop record (R, 0.93), (B, 0.85), (M, 0.81), (Q, 0.98), (S, 0.88)
@@
Refer to caption Refer to caption Refer to caption
Figure 4: Tracking the change of reconstruction loss using the B/L/W domains of the FairFace dataset.
Data Reconstruction and Generation via $T$. To assess the effectiveness of the transformation model $T$, we visualize instance reconstruction and domain variation using image datasets in Fig. 4. The first column (Original) shows the images sampled from the datasets. In the second column (Reconstruction), we display images that are generated from latent factors encoded from the images in the first column. The images in the second column closely resemble those in the first column. We showcase the reconstruction loss using the FairFace dataset in Fig. 4. Images in the last three columns are generated using the content factors that were encoded from images in the first column. These images are generated with style and sensitive factors randomly sampled from their respective prior distributions. The images in the last three columns preserve the underlying semantic information of the corresponding samples in the first column. However, their style and sensitive attributes undergo significant changes. This demonstrates that the transformation model $T$ effectively extracts latent factors and produces diverse transformations of the provided data domains. More visualization and plot results are given in Appendix F.
The Effectiveness of FDDG Across Domains in terms of Fairness and Accuracy. Comprehensive experiments showcase that FDDG consistently outperforms baselines by a considerable margin. For all tables in the main paper and Appendix, results shown in each column represent performance on the target domain, using the rest as source domains. Due to space limit, selected results for three domains of FairFace are shown in Tab. 2, but the average results are based on all domains. Complete performance for all domains of datasets refers to Appendix F. As shown in Tab. 2, for the FairFace dataset, our method has the best accuracy and fairness level for the average DG performance over all the domains. More specifically, our method has better fairness metrics (4% for DP, 2% for AUC) and comparable accuracy (0.23% better) than the best of the baselines for individual metrics. As shown in Tab. 3, for YFCC100M-FDG, our method excels in fairness metrics (8% for DP, 5% for AUC) and comparable accuracy (0.35% better) compared to the best baselines.
Ablation Studies. We conduct three ablation studies to study the robustness of FDDG on FairFace. In-depth descriptions and the pseudocodes for these studies can be found in Appendix D. More results can be found in Appendix F. (1) In FDDG w/o sf, we modify the encoder within $T$ by restricting its output to only latent content and style factors. (2) FDDG w/o $T$ skips data augmentation in synthetic domains via $T$ and results are conducted only based $f$ constrained by fair notions outlined in Defn. 1. (3) In FDDG w/o fc, the fair constraint on $f$ is not included, and we eliminate the $\mathcal{L}_{fair}$ in line 9 of Algorithm 1. We include the performance of such ablation studies in Tab. 4. The results illustrate that when data is disentangled into three factors and the model is designed accordingly, it can enhance DG performance due to covariate and dependence shifts. Generating data in synthetic domains with randomly fair dependence patterns proves to be an effective approach for ensuring fairness invariance across domains.
Table 2: Performance on for FairFace. Bold is the best and underline is the second best.
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (B, 0.91) (L, 0.48) (W, 0.49) Avg
RandAug 0.64$\pm$0.26 / 0.64$\pm$0.15 / 93.47$\pm$1.56 0.39$\pm$0.10 / 0.70$\pm$0.02 / 91.77$\pm$0.61 0.34$\pm$0.09 / 0.64$\pm$0.02 / 92.07$\pm$0.55 0.42 / 0.66 / 92.94
ERM 0.67$\pm$0.17 / 0.58$\pm$0.02 / 91.89$\pm$1.10 0.57$\pm$0.15 / 0.62$\pm$0.01 / 91.96$\pm$0.51 0.39$\pm$0.09 / 0.61$\pm$0.01 / 92.82$\pm$0.38 0.51 / 0.61 / 93.08
IRM 0.63$\pm$0.12 / 0.58$\pm$0.01 / 93.39$\pm$1.03 0.41$\pm$0.21 / 0.63$\pm$0.05 / 92.06$\pm$1.89 0.32$\pm$0.19 / 0.66$\pm$0.01 / 90.54$\pm$1.56 0.43 / 0.62 / 92.48
GDRO 0.71$\pm$0.16 / 0.57$\pm$0.02 / 89.81$\pm$1.10 0.54$\pm$0.15 / 0.62$\pm$0.01 / 91.59$\pm$0.51 0.48$\pm$0.09 / 0.60$\pm$0.01 / 92.50$\pm$0.38 0.55 / 0.60 / 92.55
Mixup 0.58$\pm$0.19 / 0.59$\pm$0.02 / 92.46$\pm$0.69 0.55$\pm$0.22 / 0.61$\pm$0.02 / 93.43$\pm$2.02 0.43$\pm$0.19 / 0.61$\pm$0.01 / 92.98$\pm$0.03 0.51 / 0.60 / 93.19
DDG 0.60$\pm$0.20 / 0.59$\pm$0.02 / 91.76$\pm$1.03 0.44$\pm$0.17 / 0.62$\pm$0.02 / 93.46$\pm$0.32 0.51$\pm$0.07 / 0.60$\pm$0.01 / 91.34$\pm$0.80 0.49 / 0.61 / 92.74
MBDG 0.60$\pm$0.15 / 0.58$\pm$0.01 / 91.29$\pm$1.41 0.56$\pm$0.09 / 0.61$\pm$0.01 / 93.49$\pm$0.97 0.30$\pm$0.04 / 0.62$\pm$0.01 / 91.05$\pm$0.53 0.50 / 0.60 / 92.71
DDG-FC 0.61$\pm$0.06 / 0.58$\pm$0.03 / 92.27$\pm$1.65 0.50$\pm$0.25 / 0.62$\pm$0.03 / 92.42$\pm$0.30 0.48$\pm$0.15 / 0.62$\pm$0.02 / 92.45$\pm$1.55 0.52 / 0.61 / 93.23
MBDG-FC 0.70$\pm$0.15 / 0.56$\pm$0.03 / 92.12$\pm$0.43 0.57$\pm$0.23 / 0.62$\pm$0.02 / 91.89$\pm$0.81 0.32$\pm$0.07 / 0.60$\pm$0.03 / 91.50$\pm$0.57 0.53 / 0.60 / 92.48
EIIL 0.88$\pm$0.07 / 0.59$\pm$0.05 / 84.75$\pm$2.16 0.49$\pm$0.07 / 0.59$\pm$0.01 / 88.39$\pm$1.25 0.46$\pm$0.05 / 0.65$\pm$0.03 / 86.53$\pm$1.02 0.64 / 0.61 / 87.78
FarconVAE 0.93$\pm$0.03 / 0.54$\pm$0.01 / 89.61$\pm$0.64 0.58$\pm$0.05 / 0.60$\pm$0.05 / 88.70$\pm$0.71 0.51$\pm$0.07 / 0.60$\pm$0.01 / 86.40$\pm$0.42 0.66 / 0.58 / 88.46
FATDM 0.93$\pm$0.03 / 0.57$\pm$0.02 / 92.20$\pm$0.36 0.51$\pm$0.16 / 0.63$\pm$0.02 / 93.33$\pm$0.20 0.46$\pm$0.05 / 0.63$\pm$0.01 / 92.56$\pm$0.31 0.67 / 0.61 / 92.54
FDDG 0.94$\pm$0.05 / 0.55$\pm$0.02 / 93.91$\pm$0.33 0.58$\pm$0.15 / 0.59$\pm$0.01 / 93.73$\pm$0.26 0.52$\pm$0.17 / 0.58$\pm$0.03 / 93.02$\pm$0.50 0.70 / 0.58 / 93.42
Table 3: Performance on YFCC100M-FDG. (bold is the best; underline is the second best).
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$

### v2

@@ 4.2 Fair Disentangled Domain Generalization
$\mathcal{L}_{fair}(\boldsymbol{\theta})=\big{|}\frac{1}{m}\sum_{(\mathbf{x}_{i},z_{i})\in\mathcal{B}}g(\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta}),z_{i})\big{|}+\big{|}\frac{1}{m}\sum_{(\mathbf{x}_{j},z_{j})\in\mathcal{B}^{\prime}}g(\hat{f}(\mathbf{x}_{j},\boldsymbol{\theta}),z_{j})\big{|}$
10:         $\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{cls}(\boldsymbol{\theta})+\lambda_{1}\cdot\mathcal{L}_{inv}(\boldsymbol{\theta})+\lambda_{2}\cdot\mathcal{L}_{fair}(\boldsymbol{\theta})$
11:         $\boldsymbol{\theta}\leftarrow\text{Adam}(\mathcal{L}(\boldsymbol{\theta}),\boldsymbol{\theta},\eta_{p})$
12:
$\lambda_{1}\leftarrow\text{max}\{[\lambda_{1}+\eta_{d}\cdot(\mathcal{L}_{inv}(\boldsymbol{\theta})-\gamma_{1})],0\}$, $\lambda_{2}\leftarrow\text{max}\{[\lambda_{2}+\eta_{d}\cdot(\mathcal{L}_{fair}(\boldsymbol{\theta})-\gamma_{2})],0\}$
@@
4.3 The FEDORA Algorithm
In practice, we propose a simple but effective algorithm, given in Algorithm 1, which is co-trained with the transformation model $T$. The detailed training process of $T$ is provided in Algorithm 2 of Appendix B. In Algorithm 1, we harness the power of $T$ to address the unconstrained dual optimization problem outlined in Sec. 4.2 through a series of primal-dual iterations.
Given a finite number of observed source domains, to enhance the generalization performance for unseen target domains, the invariant classifier $\hat{f}$ is trained by expanding the dataset with synthetic domains generated by $T$. These synthetic domains are created by introducing random sample style and random sensitive factors, hence a random sensitive attribute, resulting in an arbitrary fair dependence within such domains. As described in Fig. 2, the sensitive factor $\mathbf{a}^{s^{\prime}}$ and the style factor $\mathbf{s}^{s^{\prime}}$ are randomly sampled from their prior distributions $\mathcal{N}(0,\mathbf{I}_{a})$ and $\mathcal{N}(0,\mathbf{I}_{s})$, respectively. A sensitive attribute $z^{s^{\prime}}$ is further predicted from $\mathbf{a}^{s^{\prime}}$ through $h$. Along with the unchanged semantic factor $\mathbf{c}$ encoded by $(\mathbf{x}^{s},z^{s},y)$, they are further passed through $G$ to generate $(\mathbf{x}^{s^{\prime}},z^{s^{\prime}},y)$ with the unchanged class labels in an augmented synthetic domain. Under Assumps. 2 and 3, according to Eqs. 7 and 8, data augmented in synthetic domains are required to maintain invariance in terms of accuracy and fairness with the data in the corresponding original domains.
Specifically, in lines 15-20 of Algorithm 1, we describe the transformation procedure that takes an example $(\mathbf{x},z,y)$ as INPUT and returns an augmented example $(\mathbf{x}^{\prime},z^{\prime},y)$ from a new synthetic domain as OUTPUT. The augmented example has the same semantic factor as the input example but has different sensitive and style factors sampled from their associated prior distributions that encode a new synthetic domain. Lines 1-14 show the main training loop for FEDORA. In line 6, for each example in the minibatch $\mathcal{B}$, we apply the procedure $T$ to generate an augmented example from a new synthetic domain described above. In line 7, we consider KL-divergence as the distance metric for $d[\cdot]$. All the augmented examples are stored in the set $\mathcal{B}^{\prime}$. The Lagrangian dual loss function is defined based on $\mathcal{B}$ and $\mathcal{B}^{\prime}$ in line 10. The primal parameters $\boldsymbol{\theta}$ and the dual parameters $\lambda_{1}$ and $\lambda_{2}$ are updated in lines 11-12.
5 Analysis
With the approximation on the dual problem in Sec. 4.2, the duality gap between $P^{\star}$ in Eq. 6 and $D^{\star}_{\xi,N,\mathcal{E}_{s}}(\gamma_{1},\gamma_{2})$ in Sec. 4.2 can be explicitly bounded.
Theorem 1 (Fairness-aware Data-dependent Duality Gap).
Given $\xi>0$, assuming $\{\hat{f}(\cdot,\boldsymbol{\theta}):\boldsymbol{\theta}\in\Theta\}\subseteq\mathcal{F}$ has finite VC-dimension, with $M$ datapoints sampled from $\mathbb{P}_{XZY}$ we have
$|P^{\star}-D^{\star}_{\xi,N,\mathcal{E}_{s}}(\boldsymbol{\gamma})|\leq L||\boldsymbol{\gamma}||_{1}+\xi k(1+||\boldsymbol{\lambda}^{\star}_{p}||_{1})+O(\sqrt{\log(M)/M})$
where $\boldsymbol{\gamma}=[\gamma_{1},\gamma_{2}]^{T}$; $L$ is the Lipschitz constant of $P^{\star}(\gamma_{1},\gamma_{2})$; $k$ is a small universal constant defined in Proposition 3 of Appendix D; and $\boldsymbol{\lambda}^{\star}_{p}$ is the optimal dual variable for a perturbed version of Eq. 6.
The duality gap that arises when solving the empirical problem presented in Sec. 4.2 is minimal when the fairness-aware $T$-invariance in Defn. 3 margin $\boldsymbol{\gamma}$ is narrow, and the parametric space closely approximates $\mathcal{F}$.
Furthermore, we present the following theorem to establish an upper bound on fairness within an unseen target domain.
Theorem 2 (Fairness Upper Bound of the Unseen Target Domain).
In accordance with Defns. 1 and 8, for any domain $e\in\mathcal{E}$, the fairness dependence under instance distribution $\mathbb{P}^{e}_{XZY}$ with respect to the classifier $f\in\mathcal{F}$ is defined as:
$\rho^{e}(f)=\big{|}\mathbb{E}_{\mathbb{P}^{e}_{XZ}}g(f(X^{e}),Z^{e})\big{|}$
With observed source domains $\mathcal{E}_{s}$, the dependence at any unseen target domain $t\in\mathcal{E}\backslash\mathcal{E}_{s}$ is upper bounded. $dist[\cdot]$ is the Jensen-Shannon distance metric [50].
$\rho^{t}(f)\leq\frac{1}{|\mathcal{E}_{s}|}\sum_{s_{i}\in\mathcal{E}_{s}}\rho^{s_{i}}(f)+\sqrt{2}\min_{s_{i}\in\mathcal{E}_{s}}dist\big{[}\mathbb{P}^{t}_{XZY},\mathbb{P}^{s_{i}}_{XZY}\big{]}+\sqrt{2}\max_{s_{i},s_{j}\in\mathcal{E}_{s}}dist\big{[}\mathbb{P}^{s_{i}}_{XZY},\mathbb{P}^{s_{j}}_{XZY}\big{]}$
where $dist[\mathbb{P}_{1},\mathbb{P}_{2}]=\sqrt{\frac{1}{2}KL(\mathbb{P}_{1}||\frac{\mathbb{P}_{1}+\mathbb{P}_{2}}{2})+\frac{1}{2}KL(\mathbb{P}_{2}||\frac{\mathbb{P}_{1}+\mathbb{P}_{2}}{2})}$ is JS divergence defined based on KL divergence.
Notice that the second term in Theorem 2 becomes uncontrollable during training as it relies on the unseen target domain. Therefore, to preserve fairness across target domains, we aim to learn semantic factors that map the transformation mode $T$, ensuring that $\mathbb{P}^{s_{i}}_{C|XZY},\forall s_{i}\in\mathcal{E}_{s}$ remains invariant across source domains. Simultaneously, we strive for the classifier $f$ to achieve high fairness within the source domains. Proofs of Theorems 2 and 1 are provided in Appendix D.
@@ 6 Experiments
Datasets Domains Sensitive Attr. Labels ($s$, $\rho^{s}$), $\forall s\in\mathcal{E}_{s}$
@@
FairFace race gender age (B, 0.91), (E, 0.87), (I, 0.58),
(L, 0.48), (M, 0.87), (S, 0.39), (W, 0.49)
YFCC100M-FDG year location in-,outdoor ($d_{0}$, 0.73), ($d_{1}$, 0.84), ($d_{2}$, 0.72)
NYSF city race stop record (R, 0.93), (B, 0.85), (M, 0.81),
(Q, 0.98), (S, 0.88)
6.1 Settings
Datasets. We evaluate the performance of our FEDORA on four benchmarks. To highlight each source data and its fair dependence score $\rho^{s}$ defined in Assump. 1, we summarize the statistics in Tab. 3.
(1) ccMNIST is a domain generalization benchmark created by colorizing digits and the backgrounds of the MNIST dataset [51]. ccMNIST consists of images of handwritten digits from 0 to 9. Similar to ColoredMNIST [38], for binary classification, digits are labeled with 0 and 1 for digits from 0-4 and 5-9, respectively. ccMNIST contains 70,000 images divided into three data domains, each characterized by a different digit color (i.e., red, green, blue) and followed by a different correlation between the class label and sensitive attribute (digit background colors). (2) FairFace [52] is a dataset that contains a balanced representation of different racial groups. It includes 108,501 images from seven racial categories: Black (B), East Asian (E), Indian (I), Latino (L), Middle Eastern (M), Southeast Asian (S), and White (W). In our experiments, we set each racial group as a domain, gender as the sensitive attributes, and age ($\geq$ or $<$ 50) as the class label. (3) YFCC100M-FDG is an image dataset created by Yahoo Labs and released to the public in 2014. It is randomly selected from the YFCC100M [53] dataset with a total of 90,000 images. For domain variations, YFCC100M-FDG is divided into three domains. Each contains 30,000 images from different year ranges, before 1999 ($d_{0}$), 2000 to 2009 ($d_{1}$), and 2010 to 2014 ($d_{2}$). The outdoor or indoor tag is used as the binary class label for each image. Latitude and longitude coordinates, representing where images were taken, are translated into different continents. The North American or non-North American continent is the sensitive attribute (related to spatial disparity). (4) NYSF [54] is a real-world dataset on policing in New York City in 2011. It documents whether a pedestrian who was stopped on suspicion of weapon possession would, in fact, possess a weapon. NYSF consists of records collected in five different regions: Manhattan (M), Brooklyn (B), Queens (Q), Bronx (R), and Staten (S). We use regions as different domains. This data had a pronounced racial bias against African Americans, so we consider race (black or non-black) as the sensitive attribute.
Baselines. We compare the performance of FEDORA with 19 baseline methods that fall into two main categories: (1) 12 state-of-the-art domain generalizations methods, specifically designed to address covariate shifts: ColorJitter, ERM [37], IRM [38], GDRO [55], Mixup [56], MLDG [41], CORAL [57], MMD [40], DANN [58], CDANN [59], DDG [17], and MBDG [15], where ColorJitter is a naive function in PyTorch that randomly changes the brightness, contrast, saturation and hue of images; and (2) 7 state-of-the-art fairness-aware domain generalizations methods, specifically designed to address either covariate or dependence shifts: DDG-FC, MBDG-FC, EIIL [20], FarconVAE [19], FCR [60], FTCS [11], and FATDM [21], where DDG-FC and MBDG-FC are two baselines that built upon DDG [17] and MBDG [15], respectively by straightforwardly adding fairness constraints defined in Defn. 1 to the loss functions of the original models. Due to space limits, we present results for 14 baselines in the main paper. Comprehensive results for all baselines can be found in Appendix E.
Evaluation metrics. Three metrics are used for evaluation. Two of them are for fairness quantification, Demographic Parity ($DP$) [46] and the Area Under the ROC Curve ($AUC_{fair}$) between predictions of sensitive subgroups [61]. Notice that the $AUC_{fair}$ is not the same as the one commonly used in classification based on TPR and FPR. The intuition behind this $AUC_{fair}$ is based on the nonparametric Mann-Whitney U test, in which a fair condition is defined as the classifier’s prediction probability of a randomly selected sample $\mathbf{x}_{-1}$ from one sensitive subgroup being greater than a randomly selected sample $\mathbf{x}_{1}$ from the other sensitive subgroup is equal to the probability of $\mathbf{x}_{1}$ being greater than $\mathbf{x}_{-1}$ [5, 62]. A value of $DP$ closer to $1$ indicates fairness, and $0.5$ of $AUC_{fair}$ represents zero bias effect on predictions.
Model selection. The model selection in domain generalization is intrinsically a learning problem, followed by [15], we use leave-one-domain-out validation criteria, which is one of the three selection methods stated in [63]. Specifically, we evaluate FEDORA on the held-out source domain and average the performance of $|\mathcal{E}_{s}|-1$ domains over the held-out one.
We defer a detailed description of the experimental settings (including evaluation metrics, architectures, and hyperparameter search) in Appendix B and additional results in Appendix E.
@@ 6.1 Settings
Figure 4: Example results of generating images using latent factors encoded from three images.
Figure 5: Ablation study on FairFace. Averaged results are plotted across all domains.
Figure 6: Results of accuracy-fairness tradeoff on Fairface (left) and YFCC100M-FDG (right) sweeping over a range of $\lambda_{2}$.
Table 4: Performance on FairFace. (bold is the best; underline is the second best).
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
Methods (B, 0.91) (W, 0.49) (L, 0.48) Avg
ColorJitter 0.64$\pm$0.26 / 0.64$\pm$0.15 / 93.47$\pm$1.56 0.34$\pm$0.09 / 0.64$\pm$0.02 / 92.07$\pm$0.55 0.39$\pm$0.10 / 0.70$\pm$0.02 / 91.77$\pm$0.61 0.42 / 0.66 / 92.94
ERM 0.67$\pm$0.17 / 0.58$\pm$0.02 / 91.89$\pm$1.10 0.39$\pm$0.09 / 0.61$\pm$0.01 / 92.82$\pm$0.38 0.57$\pm$0.15 / 0.62$\pm$0.01 / 91.96$\pm$0.51 0.51 / 0.61 / 93.08
IRM 0.63$\pm$0.12 / 0.58$\pm$0.01 / 93.39$\pm$1.03 0.32$\pm$0.19 / 0.66$\pm$0.01 / 90.54$\pm$1.56 0.41$\pm$.021 / 0.63$\pm$0.05 / 92.06$\pm$1.89 0.43 / 0.62 / 92.48
GDRO 0.71$\pm$0.16 / 0.57$\pm$0.02 / 89.81$\pm$1.10 0.48$\pm$0.09 / 0.60$\pm$0.01 / 92.50$\pm$0.38 0.54$\pm$0.15 / 0.62$\pm$0.01 / 91.59$\pm$0.51 0.55 / 0.60 / 92.55
Mixup 0.58$\pm$0.19 / 0.59$\pm$0.02 / 92.46$\pm$0.69 0.43$\pm$0.19 / 0.61$\pm$0.01 / 92.98$\pm$0.03 0.55$\pm$0.22 / 0.61$\pm$0.02 / 93.43$\pm$2.02 0.51 / 0.60 / 93.19
DDG 0.60$\pm$0.20 / 0.59$\pm$0.02 / 91.76$\pm$1.03 0.51$\pm$0.07 / 0.60$\pm$0.01 / 91.34$\pm$0.80 0.44$\pm$0.17 / 0.62$\pm$0.02 / 93.46$\pm$0.32 0.49 / 0.61 / 92.74
MBDG 0.60$\pm$0.15 / 0.58$\pm$0.01 / 91.29$\pm$1.41 0.30$\pm$0.04 / 0.62$\pm$0.01 / 91.05$\pm$0.53 0.56$\pm$0.09 / 0.61$\pm$0.01 / 93.49$\pm$0.97 0.50 / 0.60 / 92.71
DDG-FC 0.61$\pm$0.06 / 0.58$\pm$0.03 / 92.27$\pm$1.65 0.48$\pm$0.15 / 0.62$\pm$0.02 / 92.45$\pm$1.55 0.50$\pm$0.25 / 0.62$\pm$0.03 / 92.42$\pm$0.30 0.52 / 0.61 / 93.23
MBDG-FC 0.70$\pm$0.15 / 0.56$\pm$0.03 / 92.12$\pm$0.43 0.32$\pm$0.07 / 0.60$\pm$0.03 / 91.50$\pm$0.57 0.57$\pm$0.23 / 0.62$\pm$0.02 / 91.89$\pm$0.81 0.53 / 0.60 / 92.48
EIIL 0.88$\pm$0.07 / 0.59$\pm$0.05 / 84.75$\pm$2.16 0.46$\pm$0.05 / 0.65$\pm$0.03 / 86.53$\pm$1.02 0.49$\pm$0.07 / 0.59$\pm$0.01 / 88.39$\pm$1.25 0.64 / 0.61 / 87.78
FarconVAE 0.93$\pm$0.03 / 0.54$\pm$0.01 / 89.61$\pm$0.64 0.51$\pm$0.07 / 0.60$\pm$0.01 / 86.40$\pm$0.42 0.58$\pm$0.05 / 0.60$\pm$0.05 / 88.70$\pm$0.71 0.66 / 0.58 / 88.46
FCR 0.81$\pm$0.05 / 0.59$\pm$0.02 / 79.66$\pm$0.25 0.39$\pm$0.06 / 0.63$\pm$0.02 / 82.33$\pm$0.89 0.38$\pm$0.12 / 0.66$\pm$0.02 / 85.22$\pm$2.33 0.54 / 0.63 / 83.68
FTCS 0.75$\pm$0.10 / 0.60$\pm$0.02 / 80.00$\pm$0.20 0.40$\pm$0.06 / 0.60$\pm$0.02 / 79.66$\pm$1.05 0.42$\pm$0.23 / 0.65$\pm$0.03 / 79.64$\pm$1.00 0.57 / 0.64 / 80.91
FATDM 0.93$\pm$0.03 / 0.57$\pm$0.02 / 92.20$\pm$0.36 0.46$\pm$0.05 / 0.63$\pm$0.01 / 92.56$\pm$0.31 0.51$\pm$0.16 / 0.63$\pm$0.02 / 93.33$\pm$0.20 0.67 / 0.61 / 92.54
FEDORA 0.94$\pm$0.05 / 0.55$\pm$0.02 / 93.91$\pm$0.33 0.52$\pm$0.17 / 0.58$\pm$0.03 / 93.02$\pm$0.50 0.58$\pm$0.15 / 0.59$\pm$0.01 / 93.73$\pm$0.26 0.70 / 0.58 / 93.42
Table 5: Performance on YFCC100M-FDG. (Bold is the best; underline is the second best.)
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$